OpenStudy (anonymous):

If x+y =10 and x-y=11, what is the value of x^2-y^2?

OpenStudy (anonymous):

110

OpenStudy (anonymous):

the best thing to do is to line up both equation and for this question we add so x+y=10 x-y=11 (x-x)+(y- -y)=(10-11) 0+2y= -1 NB: when you have two minus sign side by side you add so when 2y= -1 y=-0.5 now that we have one unknown you just put 0.5 were y is in the first equation x+0.5=10 so x=9.5 so 9.5^2-0.5^2 = 90.25-0.25= 90